OpenStudy (anonymous):

(d^2y/dx^2)+(9y)=0 y(0)=3 y'(0)=6 HOW DO YOU DO SECOND ORDER INITIAL VALUE JUST TELL ME DONT DO IT

OpenStudy (anonymous):

is this calculus based?

OpenStudy (anonymous):

diffy Q

OpenStudy (anonymous):

dont know what that means,was just thinking it my be implicit derivates

OpenStudy (anonymous):

Differential Equations

OpenStudy (anonymous):

yeah! thats implicit differentiation then

OpenStudy (amistre64):

y'' + 9y = 0

OpenStudy (amistre64):

how to explain a characteristic equation ....

OpenStudy (amistre64):

assume that e^(rx) will work y = e^(rx) y' = r e^(rx) y'' = r^2 e^(rx); and sub it int r^2 e^(rx) + 9e^(rx) = 0 factor out the es e^(rx) (r^2 + 9) = 0 since e^rx is never zero; we solve for the quadratic

OpenStudy (amistre64):

r^2 + 9 = 0 r^2 = -9 r = 0+3i after that it gets funky :) but it simplifies to: \[y=e^{0x}(c_1cos(3x)+c_2sin(3x))\] \[y=c_1cos(3x)+c_2sin(3x)\]